MaxCDN
All Systems Operational
Amsterdam
Atlanta
Chicago
Dallas
Frankfurt
Los Angeles
London
New York
Seattle
San Francisco
Singapore
Virginia
Video On Demand
Live Streaming
API
Control Panel
CDN Reports
Past Incidents
Apr 20, 2014

No incidents reported today.

Apr 19, 2014

No incidents reported.

Apr 18, 2014

No incidents reported.

Apr 17, 2014
Completed - The scheduled maintenance has been completed.
Apr 17, 15:49 PDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Apr 17, 10:00 PDT
Scheduled - We will be conducting maintenance on our devices at LAX on Thursday, April 17th, from 10am-6pm PST. During this time, traffic will be re-routed to nearby locations.
Apr 14, 13:46 PDT
Apr 16, 2014
Completed - The scheduled maintenance has been completed.
Apr 16, 12:30 PDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Apr 16, 12:00 PDT
Scheduled - We have scheduled emergency maintenance at 12:00pm PST to increase network capacity. Traffic will re-route to nearby locations, however Push file uploads will be disabled for up to 15 minutes.
Apr 16, 11:19 PDT
Apr 15, 2014
Resolved - All customer SSL certificates that were purchased through our control panel have been rekeyed. This issue is now resolved.
Apr 15, 20:17 PDT
Identified - We are currently in the process of rekeying all SSL certs that were purchased through our control panel. Fifty percent of customer certificates have already been completed. We are expecting to fully complete the rekey process by Tuesday, April 15th at 8PM PST. Should you have any questions or concerns, please contact our support team. We appreciate your patients regarding this matter.
Apr 14, 19:41 PDT
Resolved - The issue preventing DNS Entries from being created and updated has been identified. Bad DNS Records were causing a conflict and preventing proper DNS updates. The bad records were removed and DNS has synchronized properly.

All Zones which DNS was not resolving are not resolving properly.

We apologize for the inconvenience this have caused.
Apr 15, 04:13 PDT
Investigating - We are currently investigating an issue that prevents DNS Entries for newly provisioned zones from being created. Meanwhile, any newly created zone will not be resolving.

We will have an update soon.
Apr 15, 03:00 PDT
Apr 14, 2014
Resolved - This issue is now resolved. We apologize for this inconvenience.
Apr 14, 00:38 PDT
Investigating - The Control Panel and Order Form SSL Certificates are currently returning a "Certificate Revoked" Error. We are working on resolving this issue right now.

Please note that this issue does not effect clients' SSL Certificates.
Apr 13, 23:34 PDT
Apr 12, 2014

No incidents reported.

Apr 11, 2014

No incidents reported.

Apr 10, 2014
Around 9:40 PM PST we experienced a brief service disruption on Edge 21 San Francisco. Our System Engineers started investigating immediately. The issue was caused by an unforeseen traffic spike which caused an overload in nginx. The System Engineers implemented a permanent solution and the issue expected to not reoccur.

We apologize for the inconvenience.
Apr 10, 00:35 PDT
Apr 9, 2014
Resolved - Synchronization is working optimally, this issue has been resolved.
Apr 9, 13:54 PDT
Monitoring - Some customers may be experiencing delays in the synchronization of their VOD files. We are monitoring systems as the synchronization process empties its queues. This delay is expected to clear up shortly.
Apr 9, 01:52 PDT
Resolved - RTMP and Pseudo service have been restored.
Apr 9, 01:42 PDT
Investigating - We are currently experiencing issues with our Pseudo streaming and RTMP modules. We are currently investigating the issue.
Apr 8, 16:42 PDT
Apr 8, 2014
Completed - Maintenance completed
Apr 8, 18:08 PDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Apr 8, 18:00 PDT
Scheduled - Our Network Operations Team is currently working on upgrading our services. During this upgrade, our Order Form, Control Panel, and API systems will be briefly unavailable. The estimated down time will be less than 10 minutes for all of our systems. We appreciate your patience during this upgrade.
Apr 8, 12:52 PDT
Resolved - We have completed patching all vulnerable services using OpenSSL on the MaxCDN network.
Apr 8, 16:10 PDT
Identified - In response to the recent disclosure of the "Heartbleed" security vulnerability in OpenSSL, CVE-2014-0160 , we are patching all vulnerable service and assessing our general security posture across all PoPs. We have completed patching of 2/3 of our infrastructure and the remaining 1/3 is currently being patched.
Apr 8, 14:48 PDT
Resolved - This incident has been resolved.
Apr 8, 13:05 PDT
Monitoring - We have identified an issue with one of our clusters that caused higher than usual load. This has been rectified, and we are now monitoring the affected locations.
Apr 8, 12:39 PDT
Apr 7, 2014

No incidents reported.

Apr 6, 2014

No incidents reported.